OpenStudy (anonymous):

Use a calculator to approximate each to the nearest thousandth. I'm not sure how to type it into my calculator...

OpenStudy (anonymous):

\[\log _{4}3.2\]

OpenStudy (anonymous):

you do this:\[\frac{ \log_{10}(3.2) }{ \log_{10}4 }\] works with ln button too

OpenStudy (anonymous):

okay thank you!!! by the way do you know how to solve log equations? such as \[10^{n}=47\]

OpenStudy (anonymous):

yes \[b^x\] is the inverse operation of\[\log_b (x)\] therefore \[\log_b(b^x) = x\] so taking log base 10 of both sides: \[10^n = 47\]\[\log_{10}(10^n) = \log_{10} (47)\]\[n = \log47\]

OpenStudy (anonymous):

Awesome!!! Thank you!!!

OpenStudy (anonymous):

glad i could help :)